There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Saws Logigng Co. The company paid an avarage of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
Answer by Lacey020991(49)

x = douglas firs
y = ponderosa pine trees
.
x + y = 850
300x + 225y = 217,500
.
Solve by eliminating the y's
.
1) Multiply the top equation by -225
-225x + -225y = -191,250
300x + 225y = 217,500
-------------------------
Your y's will cancel when you add everything
.
75x = 26,250
x = 350
.
2) plug your x-value back into the top equation
350 + y = 850
y = 500
.
.
===================================================
You end up with 350 douglass firs and 500 ponderosa pine trees.
.
If you need further explanation, just let me know.